OpenStudy (anonymous):

please help, this is my last question of the quarter-Thank you!! A person invested $42,000 in three funds paying 5%, 7%, and 9% simple interest. The total annual interest from these investments was $2,600. The amount of money invested at 5% was $200 less than the amount invested at 7% and 9% combined. How much was invested in each fund? Check your solution

OpenStudy (rosey):

u should b a fan of me :)

OpenStudy (anonymous):

Solve the following simultaneous equations for x, y and z:\[\frac{5}{100} x+\frac{7}{100} y+\frac{9}{100} z=2600,x\text{=}(y+z)-200,x+y+z=42000\]\[\{x=20900,y=17200,z=3900\} \]\[.05*20900+.07*17200+.09*3900=2600 \]